OpenStudy (anonymous):

fing the value of p. 4/5 p =-40

OpenStudy (mayankdevnani):

p=-40*5/4=-50

OpenStudy (mayankdevnani):

@sbanuelos ok

OpenStudy (mayankdevnani):

is this incorrect?

OpenStudy (anonymous):

yes it can only ecuall ( 32 ) (-32) or (50 ) (-50)

OpenStudy (mayankdevnani):

-50

OpenStudy (mayankdevnani):

@ajprincess am i right?

OpenStudy (anonymous):

i got -32

OpenStudy (mayankdevnani):

how?

OpenStudy (mayankdevnani):

4/5*-32=-31.2

OpenStudy (mayankdevnani):

ok

OpenStudy (mayankdevnani):

4/5*-50=-40

OpenStudy (mayankdevnani):

got it

OpenStudy (mayankdevnani):

\[\frac{4}{5}\times p=-40\] \[p=-40\times \frac{5}{4}\] \[p=-10\times5\] \[p=-50\]

OpenStudy (mayankdevnani):

ok @sbanuelos